OpenStudy (anonymous):

solve by factoring x^3-7x^2-60x=0

jhonyy9 (jhonyy9):

like first step factorize out x and will get x(x^2 -7x -60) =0 - so now you need try getting two numbers so when assumed them will get -7 and when multiplie them will get -60 ,so these are for you can factorizing completely the x^2 -7x -60 - so than these two numbers will be -12 and +5 because -12 +5 = -7 and -12 *5= - 60 hope so much that now using this help you will can solving right this your exercise